OpenStudy (anonymous):

-16x^2 = 160x + 144 = 0

OpenStudy (anonymous):

divide by 16 - x^2 = 10x + 9 = 0 x^2 + 10x + 9 = 0 (x+9)(x+1) = 0 x = -1 , -9

OpenStudy (anonymous):

\[\huge 16x^2+160x + 144 = 0\] \[\huge x^2+10x + 9 = 0\] \[\huge x_1=\frac{10 + \sqrt{100-36}}{-2}\rightarrow -5-4=-9\] \[\huge x_2=\frac{10 - \sqrt{100-36}}{-2}\rightarrow -5+4 = -1\]

OpenStudy (anonymous):

scale all math by 80%

OpenStudy (anonymous):

\[\frac{-16 x^2+160 x+144}{16}=0 \]Solve the following for x:\[-x^2+10 x+9=0 \]\[\left\{x=5-\sqrt{34},x=5+\sqrt{34}\right\} \]

OpenStudy (anonymous):

\[\huge -16x^2 = 160x + 144 = 0\] not \[\huge -16x^2 - 160x + 144 = 0\]

OpenStudy (anonymous):

or \[\huge -16x^2+160x + 144 = 0\]

OpenStudy (anonymous):

Chose + as the sign of 160x because the + character on a standard keyboard is a shifted = .

OpenStudy (anonymous):

good eye

OpenStudy (anonymous):

I make mistakes myself when not holding down the shift key long enough for it to function.